Two geneticists (we’ll call them “Brad” and “Angelina”) attempt to estimate the frequency of a recessive allele in a population of beach slugs.  Brad observes that 1% of the population has the recessive phenotype; she therefore estimates the allele frequency as q = √(0.01) = 0.10.  Angelina, using a molecular test, determines that 71% of the population is homozygous for the dominant allele, and 28% is heterozygous.  Using Angelina's data, what would be the best estimate of the frequency of the recessive allele?  Why do you think this differs from Brad estimate?
Answer #1

Say frequency of the dominant allele be p.

Frequency of the homozygous dominant genotype = p​​​​​​2 = 71% = 71/100 = 0.71

Frequency of the heterozygous genotype = 2pq = 28% = 28/100 = 0.28

Since p​​​​​​2​​​​​ = 0.71, so p = 0.71 = 0.84

2pq = 0.28, or 2 x 0.84 x q = 0.28, or q = 0.28/(2 x 0.84) = 0.16

So by Angelina's data, frequency of the recessive allele q = 0.16

This value differs from Brad's estimate possibly because Brad did not take into account the recessive alleles present in heterozygotes. Brad calculated the value only from recessive homozygotes.